OpenStudy (anonymous):

Find the area of a regular pentagon with a side length of 14 and an apothem of 22. medal an fan

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mwhit2101 @sarah786 @Warriorz13 @CrazyCountryGirl @lovelyr0se @ikatouni

OpenStudy (anonymous):

@mathslover

OpenStudy (warriorz13):

Here hun try this http://www.mathopenref.com/apothem.html

OpenStudy (anonymous):

what do i do with this?

OpenStudy (warriorz13):

IDK I think it kinda like a calculator

OpenStudy (anonymous):

humm ok

OpenStudy (anonymous):

what do i do for n

OpenStudy (anonymous):

@Warriorz13

OpenStudy (warriorz13):

Let n= the number of sides of a regular polygon s=the length of each side. A= the area of the n-gon Divide the n-gon into n congruent isosceles triangles. We'll do it here for a 4-gon, where each side is 6cm. Reply Using Drawing Now divide the iso triangles into congruent right triangles. 3 3 Reply Using Drawing Figure out the angle at the middle of each right triangle. 3 x Reply Using Drawing x = 360/(n/2) = 180/n using trig: tan(x) = (s/2)/h or h = s / 2tan(180/n) the area of one of the right triangles is thus a= (1/2)(s/2)(s / (2tan(180/n))) a= s^2 / (8tan(180/n)) and the area of the n-gon is 2*n*a, so A= s^2 / (4*tan(180/n)) I would suggest working through it with the square (n=4, s=6cm) if you are still having difficulty. Good Luck

OpenStudy (anonymous):

750?